OpenStudy (mitunaampora):

PLEASE HELP ME 9 breeds of dogs with the average weights of 66,22,6.6,18,71,88,40,64,115 have the average weight together of 54.5 not much would the 10th dog have to weight for the average weight to be 250

OpenStudy (mitunaampora):

also if you could explain how you got this that would be great

Directrix (directrix):

Is this >> 6.6 meant to be 6.6 for the weight of one of the dogs?

OpenStudy (mitunaampora):

yes

Directrix (directrix):

Let x be the weight of the 10th dog. 66+22+ 6.6+ 18 +71 +88 +40 +64 +115 + x ----------------------------------------- = 250 10

Directrix (directrix):

To get the average weight from 54.5 to 250, this tenth dog will be heavy.

OpenStudy (mitunaampora):

i would imagine so lol but im not exactly sure how to calculate it

Directrix (directrix):

490.6 + x --------- = 250 10

Directrix (directrix):

490.6 + x = 250 * 10 Solve for x, the weight of the tenth dog.

Directrix (directrix):

Would you check to see if the average weight is 250?

OpenStudy (mitunaampora):

the average weight that im suppose to get is 250

OpenStudy (mitunaampora):

i think i figured it out by solving for x i got 2009.4

Directrix (directrix):

Yes. Well, the tenth dog will weigh over a ton.

OpenStudy (mitunaampora):

and i realize thats a lot but thank you SO much your a life saverrrrr

Directrix (directrix):

>> i got 2009.4 That is what I got, too.

Directrix (directrix):

You are welcome.